Youssef owns 1 share of stock A and 1 share of stock B. In 1 year from today, the total value of his holdings is expected to be 191.59 dollars. Stock A is currently priced at 88.96 dollars, has an expected return of 17.23 percent, and is expected to pay a dividend of 6.67 dollars in 1 year from today. Stock B has an expected return of 16.61 percent and is expected to pay a dividend of 3.56 dollars in 1 year from today. What is the price of stock B today?
Answer #1

Price of A 1 year from now=Price now*(1+return)-D1=88.96*(1+17.23%)-6.67=97.617808

Price of B 1 year from now=191.59-97.617808=93.972192

Price of B today=(93.972192+3.56)/(1+16.61%)=83.6396